OpenStudy (anonymous):

5. The table shows the height of a plant over time. Which type of function best models the data? Write an equation to model the data. quadratic; y = 18x2 + 85 quadratic; y = 85x2 + 18 linear; y = 18x + 85 exponential; y = 85 • 18x

OpenStudy (anonymous):

ehm, how abou the table?

OpenStudy (anonymous):

let me take a picture of it.

OpenStudy (anonymous):

we are all impatient brute math nerdisitos! lol

OpenStudy (anonymous):

Sorry my computter being a ......... So im having to send it on my phone!

OpenStudy (anonymous):

Year Height (cm) 0 85 1 103 2 121 3 139 4 157

OpenStudy (anonymous):

you could help us a lil bit out there x y 0 85 1 103 2 121 3 139 4 157

OpenStudy (anonymous):

of course

OpenStudy (anonymous):

but that is not how it looks, I had to think there... math people do not think.. ;P

OpenStudy (anonymous):

so, when x=0, y=85 among the options, which one says that? if you put x=0 in B, what is y?

OpenStudy (anonymous):

lol yes haha!

OpenStudy (anonymous):

85 ??

OpenStudy (anonymous):

in the quadratic option!

OpenStudy (anonymous):

it finally let me send it.

OpenStudy (anonymous):

im not that good with algebra js

OpenStudy (anonymous):

ok.. too late.. damage to the brain has happened!!

OpenStudy (anonymous):

huhh

OpenStudy (anonymous):

in the one that says \{y=85x^2+18\) if x=0, what is y?

OpenStudy (anonymous):

in the one that says \(y=85x^2+18\) if x=0, what is y?

OpenStudy (anonymous):

plug in x=0

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

you get y=18 right?

OpenStudy (anonymous):

oh lol yes

OpenStudy (anonymous):

so that cannot be your answer! we strike off option B :)

OpenStudy (anonymous):

There's the last question.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then check with the last one \(y=85\times18x\) if x=0, what is y?

OpenStudy (anonymous):

ummm

OpenStudy (anonymous):

ooh \(y=85\times18^x=85\times18^0=85(1)=85\) so, we are good. lets put in x=1 in A, C and D options

OpenStudy (anonymous):

Gotcha!

OpenStudy (anonymous):

and,?

OpenStudy (anonymous):

its C ?

OpenStudy (anonymous):

right on

OpenStudy (anonymous):

really ?

OpenStudy (anonymous):

now for 3 & 4 ;D